Show answers

Consider a market in which there are 9 identical firms. Marginal cost of each firm is given by MCi= 2qi, and there are no fixed costs. Market demand is given by Qd= 90- 3P.

27) Refer to Scenario 2. Assume perfect competition, so each firm is a price taker; then at market equilibrium, P= $______; Q= ______; and qi= ______.

28) Refer to Scenario 2. Assume perfect competition, ...; then at market equilibrium, each firm makes profits= $______; and total market consumers' surplus CS= $______.

29) Refer to Scenario 2. Assume perfect competition, ...; then at market equilibrium, total market producers' surplus PS= $______, and total market welfare (CS+PS) = $______.

30) Refer to Scenario 2. Suppose all the firms in the market decide to collude and act as a monopolist. Then the monopolist will charge P = $______; and produce Q= ______.

31) Refer to Scenario 2. Suppose all the firms in the market decide to collude and act as a monopolist. Then the monopolist will make profits $= ______, and consumer's surplus = $______.

32) Refer to Scenario 2. Suppose all the firms in the market decide to collude and act as a monopolist. Is each firm better-off? Does each firm make higher profits $= ______? What is deadweight loss DWL= $______.

33) Refer to Scenario 2. Suppose 5 of the existing firms discover a new technique of production that is cheaper than the old technique and they decide to merge. Consumers cannot tell the difference between the product produced by the old or new technique. The MC of the new technique is $5 and there are no fixed costs. The 5 merged firms now behave as the price leader, while the other remaining 4 firms behave as price takers. When the leader maximizes profits, PL= $______; and QL= ______. 34) Refer to Scenario 2. Suppose 5 firms discover ... , then the profits made by each one of the merged firms is $______ 35) Refer to Scenario 2. Suppose 5 firms discover ... , then the quantity produced by each one of the followers is qi= ______; and the profits made by each one of the followers is $______. 36) Refer to Scenario 2. Suppose 5 firms discover ... , then (in comparison with perfect competition) total quantity produced in the market has increased/decreased; and price has increased/decreased. 37) Refer to Scenario 2. Suppose 5 firms discover... ; then (in comparison with monopoly), do the 5 merged firms prefer monopoly or merger? What happened to profits of each of the merged firms? 38) Refer to Scenario 2. Suppose 5 firms discover... ; then (in comparison with competition), do the 5 merged firms prefer competition or merger? What happened to profits of each of the merged firms?
Answer #1

Answer: Qd = 90-P ; MCi = 2qi. Each from ss cuare gi = f Market supply a At Equilibeth Mortent 38 = Egi = 9P = 4.5P 2 (27) At

(29.) Producers Surplus (PS) - 1 x 54 x12 2 = $ 324 Total Market Welfare (cs + PS) = $486+ $324 = $810 Thus, PS = $324 cS+PS

(32) Profits; - Profits - 168.75 49 (+ = $18.75 Dead Weight Loss = 0.5X (54 - 11.25) (26.25-12) = 0.5X 42.75 X 14:25 = $ 304.